OpenStudy (anonymous):

find all solutions by factoring V2+10v+24=0

OpenStudy (anonymous):

24 = 6 * 4 (v+6) (v + 4)

OpenStudy (anonymous):

Find two numbers that multiply to give you 24 and add to give your 10: (v+6)(v+4)

OpenStudy (anonymous):

6 and 4.. seems to easy though??

OpenStudy (anonymous):

6*4=24 and 6+4=10